(2pt) Imagine a world in which 7% of people own high-tops. If4 people are chosen, what is the probability that none own high-tops? 1. Consider the following sample. DATA SET 1: 86 2. DATA SET2: 28 55 66 77 78 1) (1pt) Find the sample standard deviation for DATA SETİ: b.) lpt) Which data set is more variable?: 3. 4-11 12-19 20-27 28-35 4 2 16 17 a.) (Ipt) Find the mean for the grouped frequency distribution b.) (1pt) Find the standard deviation for the grouped frequency distribution
Answer #1

Answer to question# 1)

p = 0.07 ( or 7%)

total people selected n = 4

we need to find the probability that none of them is high tops this implies x = 0

.

We use the formula of binomial probability here as follows:

P(X=x) = nCx * p^x * (1-p)^(n-x)

on plugging the values we get:

P(X=0) = 4C0 * 0.07^0 * 0.93^4

P(x=0)= 1 *1 * 0.7481

P(X=0) = 0.7481